OpenStudy (anonymous):

Which ordered pairs from the replacement set are solutions to the equation? y = -2x + 9 ; Replacement Set {((-1, 11), (0, 8), (1, 7), (2, 5)}

OpenStudy (raden):

test one by one all order pairs, for example : y = -2x + 9 at point (0,8) 8 = -2(0) + 9 8 = 9 , this is false try it for the other points